The weights(kg) of 11 stem B follow a normal distribution and has a mean of 50 and a standard deviation of 5 how many students have weights greater than 57?

1
Expert's answer
2022-06-13T12:02:59-0400

We have a normal distribution with "\mu=50," "\sigma=5."

Let's find the corresponding z-score for "X=57" :

"z=\frac{X-\mu}{\sigma}=\frac{57-50}{5}=1.4"


So,

"P(X>57)=P(z>1.4)=1-P(z<1.4)=1-0.9192=0.0808."


As the total number of students is 11 and the probability to have weight greater than 57 is 0.0808,

the number of students of such weight equals:

"11\cdot 0.0808=0.8888\approx1."


Answer: The number of students that have weights greater than 57 is "\approx1."